# A Role of Environmental Complexity on Representation Learning in Deep Reinforcement Learning Agents

### v1

@@ 3.5.1 A small increase in shortcut experience induces a large learning benefit
Referring back to Fig 3C, we observed a large gap in learning efficiency between agents from $p=0.1$ and $p=0.8$ environments. One might assume $p=0.2$ agents would resemble those from $p=0.1$, with closed shortcut episodes dominating their experience. Yet, their learning dynamics and representation development, particularly the initial dip in landmark sensitivity, are more qualitatively similar to $p=0.5$ or $p=0.8$ agents. Additionally, $p=0.3$ agents reach nearly the same asymptotic performance as agents in $p=0.8$, far beyond those from $p=0.1$. This suggests that even small adjustments in training conditions can significantly influence learning outcomes. This could also be interpreted as navigation skills taking a critical amount of experience to develop.
While the learning mechanisms between humans and RL agents differ, we propose a modification to the DSP to test this hypothesis. Typically, DSP participants are first instructed to walk around the perimeter of a maze multiple times while noting the objects that they come across. Afterwards, they are tasked with finding their way to these objects, and the frequency that they use shortcuts in the maze (cutting through the maze, as opposed to following the path they were initially guided on) is recorded.
During the initial phase of guided exploration, we suggest incorporating some number of example paths where participants are led through the center of the maze. Based on our findings, we hypothesize that these demonstrations should induce a nonlinear improvement in the amount of shortcuts used by participants as a function of the number $N$ of shortcut demonstrations given. In other words, as $N$ is increased, we predict that initially a substantial increase in shortcut taking will be observed. Eventually though, the marginal benefit of increasing $N$ further should taper off, with the most significant gains in shortcut usage coming from the first few additional demonstrations.
@@ 3.5.2 Dependence of representation and navigation skill development on environment
In our deep RL agents, we found that the amount of training modulated the relationship between shortcut use and landmark sensitivity for across different $p$ values. This result is highlighted in Fig 16, which depicts the relationship early and late in training. Early on, lower $p$ agents showed better representation scores, due to the previously discussed dip in score that is mitigated by exposure to the landmark. However, as training progresses, the ordering reverses with higher $p$ agents achieving higher representation scores. To interpret this observation for human navigators, we equate the training time of an agent to the amount of navigation experience a human has. Evaluating the agent is similar to running a DSP trial with a human, and the representation scores that we measure in agents can be thought of as generalizable skills that an individual has picked up (comparing landmark sensitivity score in agents to landmark knowledge in humans).
Mapping our findings from artificial agents to humans then suggests that the landmark knowledge an individual from Padua or Salt Lake City has should depend on the amount of navigational experience they have. If we were to compare individuals of less experience from both groups, then those from Salt Lake City (like our $p=0.1$ agents) should have higher landmark knowledge than those from Padua (like our $p=0.8$ agents), and the relationship should reverse when comparing individuals with greater experience. To empirically test this hypothesis, one could conduct a longitudinal study tracking individuals over time to evaluate changes in their landmark knowledge. Alternatively, a factorial experiment could be designed where individuals from different environments are subdivided based on various navigational skills, such as performance on the DSP, and landmark knowledge is compared across subgroups.
@@ 4 Discussion
Our work is strongly motivated by research into individual differences in navigation skill in humans. The goal was to apply deep RL agents as a model for behavioral studies of human navigators while also taking inspiration from navigation research to develop RL training environments and representation analysis techniques. For human navigators, we provide insight into the differences found in navigation ability based on where an individual grew up. Our findings with artificial agents suggest that engaging with landmarks in sufficiently challenging environments enhances the development of landmark sensitive representations. This is similar to the findings of individuals from Padua and Salt Lake City, where the former group generally has better landmark knowledge in an environment that is harder to navigate [1]. Importantly, our analysis predicts that differences between these groups may diminish or reverse when focusing on less experienced individuals. Separately, we also predict that introducing shortcut demonstrations into a DSP experiment might lead to non-linear improvements in shortcut use frequency based on the number of demonstrations given.
We make an important note that deep RL is an imperfect model to capture the complex learning dynamics of human or animal brains. While our insights and predictions can at best roughly translate from neural network representations to human navigational skills, they are intended to inspire testable hypotheses at a qualitative level. RL also has a history of being applied to interpret brain signals, like explaining dopaminergic neuron firing rates as a temporal difference signal [5]. In our work, we measure landmark sensitive nodes in the neural network by comparing activations (which can be thought of as similar to neuron firing rates) when a landmark is present to when it is not. This technique might also be applicable in brain recordings, for example by randomly removing an object in the DSP and looking for differential activity responses to mark cue sensitive neurons.
Another important discussion point is of best practices to analyze representations in artificial agents. Our primary approach involved recording activations while the agent performed its usual policy on the task that it is trained in. This comes with the caveat that the policy that generates the trajectories itself influences the recorded activations by affecting inputs to the network, and could lead to misattributing behavioral changes to shifts in representation learning. As demonstrated in Sec 3.4.2, with artificial agents we have the option of feeding fixed, repeatable stimuli to the agent and seeing how the network responses change through training. However, we have noticed that controlling stimuli can diminish the strength of representations, and could lead to some encodings being missed entirely. In artificial agents, since representations are optimized as part of an agent’s policy, they should be considered valid within that context. Place cells, which serve as a main motivation, have also been found to be dependent on environment and task context [7, 14]. Thus, while using prescribed trajectories is a valuable technique to weigh the influence of policy changes in representation strength, representations should also be analyzed with the policies they are developed.
@@
In our representation analysis, we began with inspiration from place cells [18], mapping out locations that individual nodes were sensitive to. One technique we highlight that runs counter to this paradigm is to analyze populations of cells as a way to understand neural encodings, rather than limiting exploration to just individual nodes. We apply k-means as one way to perform analysis on a group of nodes. In Sec 3.2.2, k-means revealed common spatial features that developed in navigation learning. When matching a heatmap to its nearest cluster, it often becomes much clearer what feature that node might help encode, and these features could be hard to glean from the individual node level. In Sec 3.4, applying k-means directly to UMAP-reduced node activations revealed the existence of a trajectory-based (or “planning-based”) representation, as opposed to position-based ones we had first searched for. The key takeaway is that looking at the population of activities or neural firing rates may be a useful exploratory tool to better understand what a network or brain is encoding. In the context of the whole population of activity, noisy individual node activations reveal themselves to be part of a larger picture of useful representations, and this perspective may be broadly valuable in analyzing neural recordings.
@@ Acknowledgements.
We thank Dr. Sarah Creem-Regehr, Dr. Elliot Smith, Dr. Frederick Adler, and Dr. Akil Narayan for helpful discussions of this work. The support and resources from the Center for High Performance Computing at the University of Utah are also gratefully acknowledged.

### v2

@@ 3.5.1 A small increase in shortcut experience induces a large learning benefit
Referring back to Fig 4C, we observed a large gap in learning efficiency between agents from $p=0.1$ and $p=0.8$ environments. One might assume $p=0.2$ agents would resemble those from $p=0.1$, with closed shortcut episodes dominating their experience. Yet, their learning dynamics and representation development, particularly the initial dip in landmark sensitivity, are more qualitatively similar to $p=0.5$ or $p=0.8$ agents. Additionally, $p=0.3$ agents reach nearly the same asymptotic performance as agents in $p=0.8$, far beyond those from $p=0.1$. This suggests that even small adjustments in training conditions can significantly influence learning outcomes.
We propose a modification to the DSP to test this hypothesis. Typically, DSP participants are first instructed to walk around the perimeter of a maze multiple times while noting the objects that they come across. Afterwards, they are tasked with finding their way to these objects, and the frequency that they use shortcuts in the maze (cutting through the maze, as opposed to following the path they were initially guided on) is recorded.
During the initial phase of guided exploration, a modified DSP task could incorporate some number of example paths where participants are led through the center of the maze. We hypothesize that these demonstrations should induce a nonlinear improvement in the amount of shortcuts used by participants as a function of the number $N$ of shortcut demonstrations given. In other words, as $N$ increases from 00, we predict that initially a substantial increase in shortcut taking will be observed. Eventually the marginal benefit of increasing $N$ further should taper off, with the largest gains in shortcut usage coming from the first few added demonstrations.
@@ 3.5.2 Dependence of representation and navigation skill development on environment
In our deep RL agents, we found that amount of training modulated the relationship between shortcut use and landmark sensitivity for different $p$ values. This result is highlighted in Fig 17, which depicts the relationship early and late in training. Early on, lower $p$ agents showed better representation scores, due to the previously discussed dip in score that is mitigated by exposure to the landmark. However, as training progresses, the ordering reverses with higher $p$ agents achieving higher representation scores. This suggests a prediction in human navigators, that amount of experience may modulate the relationship between growing up environment and landmark knowledge.
Specifically, to test this hypothesis one could conduct a factorial analysis of previously collected data, where individuals from different environments are subdivided based on navigational skill. Navigation ability would act as a surrogate measure for amount of experience. We would predict that similar to the artificial agents trained on the simulated shortcut task, better landmark knowledge would be found in individuals from an environment with stable global cues but less experience. When looking at the group with more experience however, we would expect to find better landmark knowledge in those from the more challenging navigation environment.
@@ 4 Discussion
Our work is motivated by research into individual differences in navigation skill in humans. We used deep RL agents in a toy navigation task as a model for behavioral studies of human navigators while also taking inspiration from navigation research to develop RL training environments and representation analysis techniques.
Our study of learned representations in network nodes’ activity began with inspiration from place cells [18], mapping out locations that individual nodes were sensitive to. We also found nodes that were sensitive to visual stimuli provided by the pink wall cue in the simulated maze, noting that the strength of this representation was dependent on both frequency of seeing the cue, and also skill in using it as part of navigation. While the artificial agents we developed are detached from the complexities of the real world, they provide a useful model for learning, especially because of the granular access we have to the activity patterns driving their behavior. Our findings led us to predictions about how human navigators may develop internal representations and components of navigation knowledge. Specifically, we suggest that skill or amount of experience may be a key modulating factor in what is learned from different environments, and that general trends (such as one population having better landmark knowledge than another) may obfuscate more nuanced differences in individuals. A large portion of our findings come from finding ways to quantify the strength of representations, and then looking at how these scores change over training. In general, this type of strategy may be useful to explore the depth of how representations develop over time in both neural networks and brains alike.
@@
In this work, we also develop techniques to analyze encodings in populations of node rather than individual neurons. Clustering in particular was used as a tool for preliminary exploration and provides qualitative suggestions of features that are important to agents. In Sec 3.2.2, k-means revealed common spatial features that developed in navigation learning. When matching a heatmap to its nearest cluster, it becomes much clearer what feature that node might help encode. In Sec 3.4, applying k-means directly to UMAP-reduced activity data revealed the existence of a trajectory-based representation different from the position ones we had originally set out to explore. The key takeaway is that looking at the population of activities or neural firing rates may be a useful exploratory tool to better understand what a network or brain is encoding. In the context of the whole population of activity, noisy individual node activations reveal themselves to be part of a larger picture of useful representations, and this perspective may be broadly valuable in analyzing neural recordings.
Our primary approach to collecting network activity data was to record activations while the agent performed its usual policy on a set of evaluation trials. This comes with an important caveat that changes to behavior could lead to shifts in observations input to the network, leading us to mistake behavioral changes to changes in representations. As demonstrated in Sec 3.4.2, with artificial agents we have the option of feeding fixed, repeatable stimuli to the agent and seeing how the network responses change through training. However, we have noticed that controlling stimuli can diminish the strength of representations, and could lead to some encodings being missed entirely. In artificial agents, since representations are optimized as part of an agent’s policy, they should be considered valid within that context. Place cells, which serve as a main motivation, have also been found to be dependent on environment and task context [7, 14]. Thus, using prescribed trajectories is a valuable technique to control for the influence of policy changes in representation development, but representations should also be analyzed with the policies they are developed.
There are many directions that could further this research. As discussed in Sec 3.4.2, the trajectory encoding continues to strengthen even after an agent’s policy has stabilized. It would be interesting to unpack why these changes happen, possibly by examining gradients driven by the RL algorithm. Without behavioral changes, it is somewhat surprising that network representations would continue to develop. On the idea of whether certain nodes with strong individual representations are more important to behavior discussed in Sec 3.4.3, we could test the influence of these nodes on policy by performing ablation studies of them. Finally, to further validate the idea of agents in high $p$ environments developing stronger cue sensitivity with training, one could design transfer experiments where trained agents are moved to another navigation context that requires landmark usage.
@@ Acknowledgements.
This research was partially funded by the National Science Foundation. We thank Dr. Sarah Creem-Regehr, Dr. Elliot Smith, Dr. Frederick Adler, and Dr. Akil Narayan for helpful discussions of this work. The support and resources from the Center for High Performance Computing at the University of Utah are also gratefully acknowledged.